Nesto hypermarket sells three brands (A, B, and C) of laptop computers. The monthly sales of the three brands are: Brand A = 15 units; Brand B = 9 units and Brand C = 12 units. The unit price of brand A is R.O. 300; brand B is R.O. 280 and brand C is R.O. 200. The cost to the Stores is 280 R.O. for brand A; 260 R.O. for brand B, and 180 R.O for brand C. Let

 

Q = ; P =; and C =

 

Using the above matrices, calculate the following:

 

(a). Calculate total revenue 

(b). Calculate total cost 

(c). Calculate the monthly profit.



1
Expert's answer
2021-12-03T13:05:00-0500

(a). Calculation of total revenue (TR):

"TR=(Q_A \times P_A)+(Q_B \times P_B)+(Q_C \times P_C)"

"TR=(15 \times 300)+(9 \times 280)+(12 \times 200)=9420"

(b). Calculation of total cost (TC):

"TC=(Q_A \times C_A)+(Q_B \times C_B)+(Q_C \times C_C)"

"TC=(15 \times 280)+(9 \times 260)+(12 \times 180)=8700"


(c). Calculation of monthly profit:

"Profit = TR-TC"

"Profit = 9420-8700=720"